OpenStudy (solidicecream):

If Logx (1 / 8) = - 3 / 2, then x is equal to A. - 4 B. 4 C. 1 / 4 D. 10

OpenStudy (anonymous):

no one will really help you after that dumb 3+3 question

OpenStudy (solidicecream):

:/ alright

OpenStudy (anonymous):

So, transforming the equation into exponential form.... x ^ (-3/2) = 1/8 To cancel out the exponent of x... [x^(-3/2) = 1/8]^(-2/3) x = (1/8)^(-2/3) x = 8^(2/3) x = 2^2 x = 4

OpenStudy (anonymous):

i think it is right

OpenStudy (anonymous):

http://www.signazon.com/contest/grad2014/lkkajqe5/ please vote for my design

OpenStudy (solidicecream):

I turned it in and it was :D thx a lot